### v1

# Surface decomposition method for sensitivity analysis of first-passage dynamic reliability of linear systems

Abstract
This work presents a novel surface decomposition method for the sensitivity analysis of first-passage dynamic reliability of linear systems subjected to Gaussian random excitations. The method decomposes the sensitivity of first-passage failure probability into a sum of surface integrals over the constrained component limit-state hypersurfaces. The evaluation of these surface integrals can be accomplished, owing to the availability of closed-form linear expressions of both the component limit-state functions and their sensitivities for linear systems. An importance sampling strategy is introduced to further enhance the efficiency for estimating the sum of these surface integrals. The number of function evaluations required for the reliability sensitivity analysis is typically on the order of $10^{2}$ to $10^{3}$. The approach is particularly advantageous when a large number of design parameters are considered, as the results of function evaluations can be reused across different parameters. Two numerical examples are investigated to demonstrate the effectiveness of the proposed method.
keywords:
sensitivity analysis , first-passage dynamic reliability , linear systems , surface decomposition method , importance sampling
[…]
left=1in, right=1in, top=1in, bottom=1in
1 Introduction
Reliability analysis is fundamental for assessing the safety of engineering structures due to the ubiquitous uncertainties arising in structural parameters such as material properties and damping ratios, as well as in external loadings such as wind, waves and ground excitations due to earthquakes ditlevsen1996structural , der1996structural . In structural dynamics applications, the so-called first-passage dynamic reliability problems have been extensively investigated, where the system failure is defined as the event that any critical performance measure exceeds its prescribed threshold during the stochastic loading process li2009stochastic . The first-passage dynamic reliability problems can be addressed by the classical level-crossing-process-based methods with certain assumptions chen1997study , kawano1999dynamic , which, however, is generally limited to cases only involving a single response process. To overcome these limitations, various reliability approaches have been developed, such as the surrogate-modeling-based methods wang2017equivalent , xian2024physics , kim2024dimensionality , probability-density-evolution-based methods chen2007extreme , chen2019direct , xian2021seismic , lyu2022unified and sampling-based methods au2001estimation , bansal2017evaluation , xian2024relaxation .
For first-passage dynamic reliability problems of linear systems subjected to Gaussian random excitations, several advanced sampling techniques have been developed to achieve extremely efficient and highly accurate estimates of the failure probability au2001first , misraji2020application , katafygiotis2006domain , valdebenito2021failure . The common feature of these sampling techniques lies in that they fully exploit the system linearity and excitation Gaussianity. The efficient importance sampling au2001first constructs its importance density as a weighted sum of the theoretically optimal importance densities associated with the component limit-state functions, in which the weight coefficients are proportional to the component failure probabilities. This idea was further extended in directional sampling, leading to the directional importance sampling misraji2020application . The domain decomposition method katafygiotis2006domain decomposes the first-passage failure probability into a sum of domain integrals that can be evaluated by the direct Monte Carlo simulation, and an importance sampling scheme is then introduced to estimate their total contribution. The multidomain importance sampling valdebenito2021failure also leverages the idea of domain decomposition, where the resulting domain integrals are solved by line sampling, and a similar importance sampling strategy is adopted to estimate their sum.
Reliability sensitivity analysis aims to quantify and understand the influence of small perturbations of system parameters on the probability of system failure wu1994computational , lu2008reliability , papaioannou2018reliability , lacaze2015probability , talebiyan2020sampling , which is essential in risk-informed decision making song2009system , taflanidis2011simulation and reliability-based design optimization youn2004selecting , chun2016structural . Sensitivity analysis of the first-passage dynamic reliability, which represents a challenging branch of reliability sensitivity analysis, has received increasing attention in recent years au2005reliability , chun2015parameter , jensen2015reliability , yang2022structural , xian2023reliability . Au au2005reliability presented an efficient simulation approach designed for dynamic-reliability-based sensitivity analysis involving only one or two parameters by artificially constructing an augmented reliability problem. Chun et al. chun2015parameter developed a sequential compounding method for the parametric sensitivity analysis of general system problems, and then applied this method to a first-passage dynamic reliability problem of a linear system. Jensen et al. jensen2015reliability investigated a subset-simulation-based approach to compute the sensitivities of first-excursion failure probability with respect to the distribution parameters of uncertain design variables. Yang et al. yang2022structural extended the probability density evolution method for evaluation of the first-passage dynamic reliability sensitivities of nonlinear structures with respect to deterministic design parameters. Xian and Su xian2023reliability developed an efficient approach for sensitivity analysis of the first-excursion dynamic reliability of fractionally-damped structures incorporating the explicit time-domain method, adjoint variable method and Poisson process method. Furthermore, building upon the developments confined to linear systems under Gaussian random excitations, the directional importance sampling and domain decomposition method have been extended from dynamic reliability analysis to reliability sensitivity analysis in valdebenito2021sensitivity and misraji2025first , respectively.
In essence, reliability sensitivity analysis entails the evaluation of a high-dimensional integral over the limit-state hypersurface, which is generally challenging to compute papaioannou2018reliability , particularly for first-passage dynamic reliability problems characterized by highly non-smooth and nonlinear limit-state hypersurfaces. Fortunately, in the case of linear systems subjected to Gaussian excitations, the above complex surface integrals can be evaluated with remarkable efficiency. This work presents a novel surface decomposition method to accomplish this purpose. The core concept of this method is to decompose the complex surface integral into a sum of simpler surface integrals over the constrained component limit-state hypersurfaces. The so-called constrained component hypersurface corresponds exactly to the active part of the corresponding component hypersurface that contributes to the system limit-state hypersurface. The effectiveness of this strategy lies in the ease of computing these component surface integrals, as closed-form linear expressions of both the component limit-state functions and their sensitivities are available. An importance sampling scheme is further introduced to estimate the total contribution of these component surface integrals. The sensitivity analyses of the first-passage dynamic reliability problems for an oscillator and a 20-degree-of-freedom shear-type structure are conducted to demonstrate the feasibility of the proposed method.
2 Explicit formulation of dynamic response and sensitivity for linear systems
2.1 Explicit expressions of dynamic responses
The equation of motion for a linear system can be expressed as
$\boldsymbol{M}\ddot{\boldsymbol{U}}(\boldsymbol{X},t)+\boldsymbol{C}\dot{\boldsymbol{U}}(\boldsymbol{X},t)+\boldsymbol{K}\boldsymbol{U}(\boldsymbol{X},t)=\boldsymbol{L}F(\boldsymbol{X},t)$ (1)
where $\boldsymbol{M}$, $\boldsymbol{C}$ and $\boldsymbol{K}$ are the mass, damping and stiffness matrices of the linear system, respectively; $\ddot{\boldsymbol{U}}(\boldsymbol{X},t)$, $\dot{\boldsymbol{U}}(\boldsymbol{X},t)$ and $\boldsymbol{U}(\boldsymbol{X},t)$ are the acceleration, velocity and displacement vectors of the linear system, respectively; $F(\boldsymbol{X},t)$ is the external excitation assumed to be a zero-mean Gaussian random process; $\boldsymbol{X}$ is the input random vector associated with the external excitation; and $\boldsymbol{L}$ is the orientation vector of $F(\boldsymbol{X},t)$.
Suppose that the linear system is initially at rest. Then, from the discrete perspective of the convolution integral, any response $r(\boldsymbol{X},t)$ of the linear system can be written in an explicit time-domain form as follows:
$r_{i}(\boldsymbol{X})=a^{r}_{i,1}F_{1}(\boldsymbol{X})+a^{r}_{i,2}F_{2}(\boldsymbol{X})+\cdots+a^{r}_{i,i}F_{i}(\boldsymbol{X})\quad(i=1,2,\dots,n)$ (2)
where $n$ is the number of time steps for the time-history analysis; $r_{i}(\boldsymbol{X})=r(\boldsymbol{X},t_{i})$ with $t_{i}=i\Delta t$ and $\Delta t$ being the time step; $F_{j}(\boldsymbol{X})=F(\boldsymbol{X},t_{j})$ with $t_{j}=j\Delta t\;(j=1,2,\dots,i)$; $a^{r}_{i,j}\;(j=1,2,\dots,i)$ are the coefficients with respect to $r_{i}(\boldsymbol{X})$.
The physical meaning and computational scheme of the coefficients $a^{r}_{i,j}\;(1\leq j\leq i\leq n)$ have been well discussed in su2014random . For these coefficients, only $a^{r}_{i,1}\;(1\leq i\leq n)$ need to be calculated and stored, while the other coefficients can be determined from the recursive formula $a^{r}_{i,j}=a^{r}_{i-1,j-1}\;(2\leq j\leq i\leq n)$. The coefficients $a^{r}_{i,1}\;(1\leq i\leq n)$ represent the response $r(\boldsymbol{X},t)$ of the linear system at time $t=t_{1},t_{2},\dots,t_{n}$ subjected to a unit impulse excitation applied at time $t=t_{1}$. Therefore, by performing a single response time-history analysis of the linear system shown in Eq.(1) with any type of time-domain integration methods, the coefficients $a^{r}_{i,j}\;(1\leq j\leq i\leq n)$ with respect to any response $r(\boldsymbol{X},t)$ of interest can be readily obtained.
The Gaussian random excitation $F(\boldsymbol{X},t)$ can be characterized by either its correlation function $R_{F}(t,\tau)$ or its power spectrum $S_{F}(\omega,t)$. Given the correlation function $R_{F}(t,\tau)$, the excitation at time $t=t_{i}$ can be represented, based on the orthogonal decomposition method grigoriu2006evaluation , as
$F_{i}(\boldsymbol{X})=\boldsymbol{\psi}_{i}\boldsymbol{X}\quad(i=1,2,\dots,n)$ (3)
where $\boldsymbol{X}$ is a standard Gaussian random vector with a dimension of $d=n$; and $\boldsymbol{\psi}_{i}$ is the $i$-th row vector of the matrix $\boldsymbol{\Psi}\boldsymbol{\Lambda}^{1/2}$, in which $\boldsymbol{\Psi}$ and $\boldsymbol{\Lambda}$ are respectively the orthonormal eigenvector matrix and the diagonal eigenvalue matrix of the following covariance matrix:
$\boldsymbol{\Sigma}=\begin{bmatrix}R_{F}(t_{1},t_{1})&R_{F}(t_{1},t_{2})&\cdots&R_{F}(t_{1},t_{n})\\ R_{F}(t_{2},t_{1})&R_{F}(t_{2},t_{2})&\cdots&R_{F}(t_{2},t_{n})\\ \vdots&\vdots&\ddots&\vdots\\ R_{F}(t_{n},t_{1})&R_{F}(t_{n},t_{2})&\cdots&R_{F}(t_{n},t_{n})\end{bmatrix}$ (4)
If the power spectrum $S_{F}(\omega,t)$ is available, the excitation at time $t=t_{i}$ can also be expressed in the same form as Eq.(3) utilizing the spectral representation method deodatis2025spectral . In this case, $\boldsymbol{\psi}_{i}$ is written as
$\boldsymbol{\psi}_{i}=\sqrt{2\Delta\omega}\begin{bmatrix}\sqrt{S_{F}(\omega_{1},t_{i})}\cos(\omega_{1}t_{i})\\ \sqrt{S_{F}(\omega_{2},t_{i})}\cos(\omega_{2}t_{i})\\ \vdots\\ \sqrt{S_{F}(\omega_{q},t_{i})}\cos(\omega_{q}t_{i})\\ \sqrt{S_{F}(\omega_{1},t_{i})}\sin(\omega_{1}t_{i})\\ \sqrt{S_{F}(\omega_{2},t_{i})}\sin(\omega_{2}t_{i})\\ \vdots\\ \sqrt{S_{F}(\omega_{q},t_{i})}\sin(\omega_{q}t_{i})\end{bmatrix}^{\mathrm{T}}\quad(i=1,2,\dots,n)$ (5)
where $q$ is the number of circular frequency intervals; $\Delta\omega=(\omega_{\max}-\omega_{\min})/q$ is the circular frequency interval with $\omega_{\max}$ and $\omega_{\min}$ being the upper and lower cutoff circular frequencies, respectively; and $\omega_{k}=\omega_{\min}+(k-0.5)\Delta\omega\;(k=1,2,\dots,q)$ is the circular frequency in the middle interval. Note that, when $\boldsymbol{\psi}_{i}$ is expressed in Eq.(5), the dimension of the standard Gaussian random vector $\boldsymbol{X}$ is $d=2q$.
Substitution of Eq.(3) into Eq.(2) yields a more compact explicit expression of $r(\boldsymbol{X},t)$ as follows:
$r_{i}(\boldsymbol{X})=\boldsymbol{a}_{i}^{r}\boldsymbol{X}\quad(i=1,2,\dots,n)$ (6)
where
$\boldsymbol{a}_{i}^{r}=a^{r}_{i,1}\boldsymbol{\psi}_{1}+a^{r}_{i,2}\boldsymbol{\psi}_{2}+\cdots+a^{r}_{i,i}\boldsymbol{\psi}_{i}$ (7)
is the coefficient vector with respect to $r_{i}(\boldsymbol{X})$. Eq.(6) indicates that the response of the linear system at any given time instant can be written as a linear combination of all the random variables contained in the input random vector $\boldsymbol{X}$.
It is noted that, for brevity, this study merely focuses on a single random excitation in the problem formulation. However, the extension to cases involving multiple random excitations is straightforward, with the number of response time-history analyses required for explicit formulation being equal to the number of random excitations.
2.2 Explicit expressions of response sensitivities
Suppose $\theta$ is a design parameter of the linear system shown in Eq.(1). Then, differentiating all the terms in Eq.(1) with respect to the design parameter $\theta$, one can derive the sensitivity equation for the linear system as follows:
$\boldsymbol{M}\frac{\partial\ddot{\boldsymbol{U}}(\boldsymbol{X},t)}{\partial\theta}+\boldsymbol{C}\frac{\partial\dot{\boldsymbol{U}}(\boldsymbol{X},t)}{\partial\theta}+\boldsymbol{K}\frac{\partial\boldsymbol{U}(\boldsymbol{X},t)}{\partial\theta}$ (8)
$=\frac{\partial\boldsymbol{L}}{\partial\theta}F(\boldsymbol{X},t)-\left\{\frac{\partial\boldsymbol{M}}{\partial\theta}\ddot{\boldsymbol{U}}(\boldsymbol{X},t)+\frac{\partial\boldsymbol{C}}{\partial\theta}\dot{\boldsymbol{U}}(\boldsymbol{X},t)+\frac{\partial\boldsymbol{K}}{\partial\theta}\boldsymbol{U}(\boldsymbol{X},t)\right\}$
where $\partial\boldsymbol{M}/\partial\theta$, $\partial\boldsymbol{C}/\partial\theta$, $\partial\boldsymbol{K}/\partial\theta$ and $\partial\boldsymbol{L}/\partial\theta$ denote the sensitivities of the matrices $\boldsymbol{M}$, $\boldsymbol{C}$, $\boldsymbol{K}$ and $\boldsymbol{L}$ with respect to $\theta$, respectively; and $\partial\ddot{\boldsymbol{U}}(\boldsymbol{X},t)/\partial\theta$, $\partial\dot{\boldsymbol{U}}(\boldsymbol{X},t)/\partial\theta$ and $\partial\boldsymbol{U}(\boldsymbol{X},t)/\partial\theta$ denote the sensitivities of the response vectors $\ddot{\boldsymbol{U}}(\boldsymbol{X},t)$, $\dot{\boldsymbol{U}}(\boldsymbol{X},t)$ and $\boldsymbol{U}(\boldsymbol{X},t)$ with respect to $\theta$, respectively.
Differentiation of both sides of Eq.(2) with respect to the design parameter $\theta$ yields the explicit time-domain form of the response sensitivity $\partial r(\boldsymbol{X},t)/\partial\theta$ as follows:
$\frac{\partial r_{i}(\boldsymbol{X})}{\partial\theta}=b^{r}_{i,1}F_{1}(\boldsymbol{X})+b^{r}_{i,2}F_{2}(\boldsymbol{X})+\cdots+b^{r}_{i,i}F_{i}(\boldsymbol{X})\quad(i=1,2,\dots,n)$ (9)
where $b^{r}_{i,j}=\partial a^{r}_{i,j}/\partial\theta\;(j=1,2,\dots,i)$ are the coefficients with respect to $\partial r_{i}(\boldsymbol{X})/\partial\theta$.
The coefficients $b^{r}_{i,j}\;(1\leq j\leq i\leq n)$ can be derived via direct differentiation of the coefficients $a^{r}_{i,j}\;(1\leq j\leq i\leq n)$ in Eq.(2) with respect to $\theta$. However, this requires knowledge of the analytical formulas of $a^{r}_{i,j}\;(1\leq j\leq i\leq n)$. In fact, similar to the coefficients $a^{r}_{i,j}\;(1\leq j\leq i\leq n)$, the coefficients $b^{r}_{i,j}\;(1\leq j\leq i\leq n)$ also possess inherent physical meanings, and a detailed computational scheme of these coefficients have been presented in hu2016explicit . Among these coefficients, only $b^{r}_{i,1}\;(1\leq i\leq n)$ need to be calculated and stored, while a recursive formula $b^{r}_{i,j}=b^{r}_{i-1,j-1}\;(2\leq j\leq i\leq n)$ can be adopted to determine the other coefficients. The coefficients $b^{r}_{i,1}\;(1\leq i\leq n)$ represent the response sensitivity $\partial r(\boldsymbol{X},t)/\partial\theta$ of the linear system at time $t=t_{1},t_{2},\dots,t_{n}$ driven by a unit impulse excitation applied at time $t=t_{1}$. Therefore, by conducting a single response sensitivity time-history analysis based on the sensitivity equation of the linear system presented in Eq.(8), the coefficients $b^{r}_{i,j}\;(1\leq j\leq i\leq n)$ can be easily obtained for the sensitivity of any response $\partial r(\boldsymbol{X},t)/\partial\theta$. The above response sensitivity time-history analysis can be achieved using any type of time-domain integration methods.
Similar to the derivation of Eq.(6), substitution of Eq.(3) into Eq.(9) results in the following compact explicit expression of $\partial r(\boldsymbol{X},t)/\partial\theta$ as
$\frac{\partial r_{i}(\boldsymbol{X})}{\partial\theta}=\boldsymbol{b}_{i}^{r}\boldsymbol{X}\quad(i=1,2,\dots,n)$ (10)
where
$\boldsymbol{b}_{i}^{r}=b^{r}_{i,1}\boldsymbol{\psi}_{1}+b^{r}_{i,2}\boldsymbol{\psi}_{2}+\cdots+b^{r}_{i,i}\boldsymbol{\psi}_{i}$ (11)
is the coefficient vector with respect to $\partial r_{i}(\boldsymbol{X})/\partial\theta$.
It is noted that, when $\theta$ is associated with the external excitation $F(\boldsymbol{X},t)$, for instance as a parameter defining the correlation function $R_{F}(t,\tau)$ or the power spectrum $S_{F}(\omega,t)$, the explicit expression of $\partial r(\boldsymbol{X},t)/\partial\theta$ can be obtained by direct differentiation of Eqs.(6) and (7), without performing any additional response sensitivity time-history analysis of the linear system.
Thus far, the explicit expressions for both the dynamic responses and their sensitivities have been constructed for the linear system, as shown in Eqs.(6) and (10), respectively. The availability of such concise closed-form expressions greatly facilitates the sensitivity analysis of first-passage dynamic reliability, which will be demonstrated in Section 3.
3 Surface decomposition method for sensitivity analysis of first-passage
dynamic reliability
3.1 First-passage dynamic reliability
For first-passage dynamic reliability problems of linear systems, the system failure event is typically defined as the exceedance of the prescribed threshold by any response component of interest at any time instant within a specified time duration. The system failure event can be formulated as the union of the component failure events, i.e.
$F=\bigcup_{k=1}^{m}\bigcup_{i=1}^{n}F_{ki}$ (12)
where $m$ is the number of critical response components; $n$ is the number of time steps for the given time interval; and $F_{ki}$ is the component failure event corresponding to the $k$th critical response component $s_{k}(\boldsymbol{X},t)$ at the time instant $t_{i}$, which can be defined as
$F_{ki}=\left\{c_{k}-s_{k,i}(\boldsymbol{X})\leq 0\right\}\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (13)
where $s_{k,i}(\boldsymbol{X})=s_{k}(\boldsymbol{X},t_{i})$; and $c_{k}$ is the positive threshold for the $k$th critical response component. Note that the formulation is presented for up-crossing of positive thresholds, while extension to down-crossing of negative thresholds or simultaneous consideration of both cases is straightforward.
The probability of the component failure event $F_{ki}$ is given by
$\mathbb{P}_{ki}=\int_{\mathbb{R}^{n}}\mathcal{H}(-g_{ki}(\boldsymbol{x}))f_{\boldsymbol{X}}(\boldsymbol{x})\,d\boldsymbol{x}\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (14)
where
$g_{ki}(\boldsymbol{x})=c_{k}-s_{k,i}(\boldsymbol{x})\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (15)
is the component limit-state function with respect to $F_{ki}$; $f_{\boldsymbol{X}}(\boldsymbol{x})$ is the joint probability density function (PDF) of the standard Gaussian random vector $\boldsymbol{X}$; and $\mathcal{H}(\cdot)$ denotes the Heaviside step function, with $\mathcal{H}(-g_{ki}(\boldsymbol{x}))$ taking 1 if $g_{ki}(\boldsymbol{x})\leq 0$ and 0 otherwise.
Let $r(\boldsymbol{X},t)$ be $s_{k}(\boldsymbol{X},t)$. Substituting Eq.(6) into Eq.(15), one has
$g_{ki}(\boldsymbol{x})=c_{k}-\boldsymbol{a}_{i}^{s_{k}}\boldsymbol{x}\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (16)
Since the limit-state hypersurface described by $g_{ki}(\boldsymbol{x})=c_{k}-\boldsymbol{a}_{i}^{s_{k}}\boldsymbol{x}=0$ is a hyperplane, closed-form solution to the failure probability $\mathbb{P}_{ki}$ can be obtained as
$\mathbb{P}_{ki}=\Phi(-\beta_{ki})\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (17)
where $\Phi(\cdot)$ is the standard Gaussian cumulative distribution function (CDF); and $\beta_{ki}$ is the reliability index with respect to $F_{ki}$, which can be expressed as
$\beta_{ki}=\frac{\mu_{g_{ki}}}{\sigma_{g_{ki}}}=\frac{c_{k}}{\left\|\boldsymbol{a}_{i}^{s_{k}}\right\|}\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (18)
where $\mu_{g_{ki}}$ and $\sigma_{g_{ki}}$ are the mean and standard deviation of the Gaussian random variable $g_{ki}(\boldsymbol{X})$; and $\left\|\cdot\right\|$ is the Euclidean norm.
The reliability index $\beta_{ki}$ can be interpreted geometrically as the Euclidean distance between the origin and the design point $\boldsymbol{x}_{ki}^{\star}$. The design point $\boldsymbol{x}_{ki}^{\star}$, also referred to as the most probable point, is defined as the point on the limit-state hypersurface $g_{ki}(\boldsymbol{x})=0$ that lies closest to the origin and can be expressed as
$\boldsymbol{x}_{ki}^{\star}=\beta_{ki}\boldsymbol{u}_{ki}\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (19)
where $\boldsymbol{u}_{ki}$ is the unit vector pointing from the origin to the design point, which is aligned with the normalized gradient of the limit-state function $g_{ki}(\boldsymbol{x})$ and hence perpendicular to the hyperplane $g_{ki}(\boldsymbol{x})=0$. The unit vector is given by
$\boldsymbol{u}_{ki}=-\frac{\nabla_{\boldsymbol{x}}g_{ki}(\boldsymbol{x})}{\left\|\nabla_{\boldsymbol{x}}g_{ki}(\boldsymbol{x})\right\|}=\frac{\boldsymbol{a}_{i}^{s_{k}}}{\left\|\boldsymbol{a}_{i}^{s_{k}}\right\|}\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (20)
where $\nabla_{\boldsymbol{x}}g_{ki}(\boldsymbol{x})$ denotes the gradient of $g_{ki}(\boldsymbol{x})$.
Although the probabilities of individual component failure events can be readily computed in closed form by Eqs.(17) and (18), an analytical evaluation of the system failure probability is not feasible. The probability of the system failure event $F$ can be expressed as
$\mathbb{P}=\int_{\mathbb{R}^{n}}\mathcal{H}(-G(\boldsymbol{x}))f_{\boldsymbol{X}}(\boldsymbol{x})\,d\boldsymbol{x}$ (21)
where $G(\boldsymbol{x})$ is the system limit-state function written as
$G(\boldsymbol{x})=\min_{k=1}^{m}\min_{i=1}^{n}g_{ki}(\boldsymbol{x})$ (22)
By exploiting the system linearity and leveraging the analytical solutions to the component failure probabilities, certain advanced simulation techniques, such as the efficient importance sampling au2001first and the domain decomposition method katafygiotis2006domain , were proposed for solving the probability integral in Eq.(21) at high efficiency.
3.2 Sensitivity of first-passage dynamic reliability
Differentiation of the system failure probability formulated in Eq.(21) with respect to the design parameter $\theta$ leads to the sensitivity of system failure probability as follows:
$\frac{\partial\mathbb{P}}{\partial\theta}=-\int_{\mathbb{R}^{n}}\delta(-G(\boldsymbol{x}))\frac{\partial G(\boldsymbol{x})}{\partial\theta}f_{\boldsymbol{X}}(\boldsymbol{x})\,d\boldsymbol{x}$ (23)
where $\delta(\cdot)$ denotes the Dirac delta function; and $\partial G(\boldsymbol{x})/\partial\theta$ is the sensitivity of the system limit-state function $G(\boldsymbol{x})$ with respect to $\theta$.
Considering the system limit-state function shown in Eq.(22), the term $\delta(-G(\boldsymbol{x}))$ in Eq.(23) can be expanded as
$\delta(-G(\boldsymbol{x}))=\sum_{k=1}^{m}\sum_{i=1}^{n}\delta(-g_{ki}(\boldsymbol{x}))\mathbb{I}_{ki}(\boldsymbol{x})$ (24)
where $\mathbb{I}_{ki}(\boldsymbol{x})$ denotes the indicator function that gives 1 if $g_{ki}(\boldsymbol{x})$ attains the minimum among all component limit-state functions, and 0 otherwise. The indicator function can be expressed as
$\mathbb{I}_{ki}(\boldsymbol{x})=\begin{cases}1,&\text{if }g_{ki}(\boldsymbol{x})=\min\limits_{j=1}^{m}\min\limits_{s=1}^{n}g_{js}(\boldsymbol{x}),\\[6.0pt] 0,&\text{otherwise}.\end{cases}\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (25)
The expansion presented in Eq.(24) decomposes the delta function of the system limit-state function into a sum of the delta functions of the individual component limit-state functions, with the indicator functions serving to partition the domain such that only the component reaching the minimum is active within each subdomain. It is noted that the indicator functions shown in Eq.(25) also provide an effective way to partition the system failure domain in the domain decomposition method for first-passage dynamic reliability analysis.
Differentiating the system limit-state function in Eq.(22) with respect to the design parameter $\theta$, and leveraging the indicator functions defined in Eq.(25), one can obtain the sensitivity of the system limit-state function as
$\frac{\partial G(\boldsymbol{x})}{\partial\theta}=\sum_{k=1}^{m}\sum_{i=1}^{n}\frac{\partial g_{ki}(\boldsymbol{x})}{\partial\theta}\mathbb{I}_{ki}(\boldsymbol{x})$ (26)
where $\partial g_{ki}(\boldsymbol{x})/\partial\theta$ is the sensitivity of the component limit-state function $g_{ki}(\boldsymbol{x})$ with respect to $\theta$, which can be derived from Eq.(15) as follows:
$\frac{\partial g_{ki}(\boldsymbol{x})}{\partial\theta}=-\frac{\partial s_{k,i}(\boldsymbol{x})}{\partial\theta}\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (27)
where $\partial s_{k,i}(\boldsymbol{x})/\partial\theta$ is the sensitivity of the $k$th critical response component $s_{k}(\boldsymbol{X},t)$ at the time instant $t_{i}$ with respect to $\theta$. Let $\partial r(\boldsymbol{X},t)/\partial\theta=\partial s_{k}(\boldsymbol{X},t)/\partial\theta$. Substituting Eq.(10) into Eq.(27), one gets
$\frac{\partial g_{ki}(\boldsymbol{x})}{\partial\theta}=-\boldsymbol{b}_{i}^{s_{k}}\boldsymbol{x}\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (28)
Substitution of Eqs.(24) and (26) into Eq.(23) yields the expansion for the sensitivity of system failure probability as follows:
$\frac{\partial\mathbb{P}}{\partial\theta}=\sum_{k=1}^{m}\sum_{i=1}^{n}\eta_{ki}$ (29)
where
$\eta_{ki}=-\int_{\mathbb{R}^{n}}\delta(-g_{ki}(\boldsymbol{x}))\frac{\partial g_{ki}(\boldsymbol{x})}{\partial\theta}\mathbb{I}_{ki}(\boldsymbol{x})f_{\boldsymbol{X}}(\boldsymbol{x})\,d\boldsymbol{x}\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (30)
Since each component limit-state function $g_{ki}(\boldsymbol{x})$ is continuously differentiable with a non-zero constant gradient, the $d$-dimensional volume integral in Eq.(30) can be converted into a (d1)(d-1)-dimensional surface integral as follows hormander2003analysis :
$\eta_{ki}=-\int_{g_{ki}(\boldsymbol{x})=0}\frac{1}{\left\|\nabla_{\boldsymbol{x}}g_{ki}(\boldsymbol{x})\right\|}\frac{\partial g_{ki}(\boldsymbol{x})}{\partial\theta}\mathbb{I}_{ki}(\boldsymbol{x})f_{\boldsymbol{X}}(\boldsymbol{x})\,dS(\boldsymbol{x})\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (31)
where $dS(\boldsymbol{x})$ denotes the differential element of the (d1)(d-1)-dimensional hyperplane defined by $g_{ki}(\boldsymbol{x})=0$; and $\left\|\nabla_{\boldsymbol{x}}g_{ki}(\boldsymbol{x})\right\|$ is the Euclidean norm of the gradient of the component limit-state function $g_{ki}(\boldsymbol{x})$, which can be obtained from Eq.(16) as follows:
$\left\|\nabla_{\boldsymbol{x}}g_{ki}(\boldsymbol{x})\right\|=\left\|\boldsymbol{a}_{i}^{s_{k}}\right\|\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (32)
Since the integration in Eq.(31) is performed over the (d1)(d-1)-dimensional hyperplane $g_{ki}(\boldsymbol{x})=0$, the indicator function shown in Eq.(25) can be further expressed as
[…]
Essentially, the indicator function $\mathbb{I}_{ki}(\boldsymbol{x})$ imposes additional constraints, restricting the integration to the portion of the hyperplane where all other component limit-state functions remain in their safe domains.
It should be noted that the constrained hyperplane $g_{ki}(\boldsymbol{x})\mathbb{I}_{ki}(\boldsymbol{x})=0$ corresponds exactly to the portion of the system limit-state hypersurface contributed by the corresponding component limit-state hypersurface. The union of all these constrained hyperplanes, which are mutually exclusive, constitutes the complete system limit-state hypersurface. Consequently, computing the sensitivity of the first-passage dynamic reliability amounts to integrating over a non-smooth system limit-state hypersurface. In this sense, Eqs.(29) and (31) can be interpreted as decomposing this complex surface integral into a set of simpler integrals over the constrained hyperplanes—a strategy referred to as the surface decomposition method. The basic idea of the proposed surface decomposition method is further illustrated in Figure 1 using a 2-dimensional toy problem.
Figure 1: Illustration of surface decomposition method using a 2-dimensional problem. Consider a series system limit-state function $G(\boldsymbol{x})=\min_{k=1}^{4}g_{k}(\boldsymbol{x})$. The four component limit-state functions are $g_{1}(\boldsymbol{x})=12-(-2x_{1}+x_{2})$, $g_{2}(\boldsymbol{x})=18-(-x_{1}+3x_{2})$, $g_{3}(\boldsymbol{x})=36-(6x_{1}+7x_{2})$ and $g_{4}(\boldsymbol{x})=10-(2x_{1}-x_{2})$, respectively. The system failure boundary is composed of four constrained component failure boundaries $g_{k}(\boldsymbol{x})\mathbb{I}_{k}(\boldsymbol{x})=0\;(k=1,2,3,4)$, in which the indicator function $\mathbb{I}_{k}(\boldsymbol{x})$ takes 1 if $g_{j}(\boldsymbol{x})>0,\forall j\neq k$, and 0 otherwise. The sensitivity of series system failure probability can be decomposed into four line integrals over these four constrained component failure boundaries, respectively. These line integrals will become surface integrals over the constrained component failure surfaces in higher-dimensional problems.
Thus far, substituting Eqs.(28) and (32) into Eq.(31), the surface integral $\eta_{ki}$ can be directly estimated with the Monte Carlo simulation as follows:
$\eta_{ki}\approx\frac{1}{N_{ki}}\sum_{j=1}^{N_{ki}}\frac{\boldsymbol{b}_{i}^{s_{k}}\boldsymbol{x}_{j}}{\left\|\boldsymbol{a}_{i}^{s_{k}}\right\|}\mathbb{I}_{ki}(\boldsymbol{x}_{j})\varphi(\beta_{ki})\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (34)
where $\varphi(\cdot)$ is the standard Gaussian PDF; $\beta_{ki}$ is the reliability index associated with the component limit-state function $g_{ki}(\boldsymbol{x})$, whose closed-form solution is given in Eq.(18); and $\boldsymbol{x}_{j}\;(j=1,2,\dots,N_{ki})$ are samples generated from the PDF $f_{\boldsymbol{X}}(\boldsymbol{x})$ lying on the hyperplane $g_{ki}(\boldsymbol{x})=0$, with $N_{ki}$ being the number of samples.
Direct simulation of sample points on the hyperplane $g_{ki}(\boldsymbol{x})=0$ can be accomplished as follows: i) generate a sample point $\boldsymbol{x}$ based on the PDF $f_{\boldsymbol{X}}(\boldsymbol{x})$; ii) decompose $\boldsymbol{x}$ into a component that aligns with the unit vector $\boldsymbol{u}_{ki}$, $\boldsymbol{x}_{\parallel}=(\boldsymbol{x}^{\mathrm{T}}\boldsymbol{u}_{ki})\boldsymbol{u}_{ki}$, and a component that is perpendicular to $\boldsymbol{x}_{\parallel}$, $\boldsymbol{x}_{\perp}=\boldsymbol{x}-\boldsymbol{x}_{\parallel}$; iii) construct ${\boldsymbol{x}}^{\prime}=\beta_{ki}\boldsymbol{u}_{ki}+\boldsymbol{x}_{\perp}$, which represents a sample point lying independently on the hyperplane $g_{ki}(\boldsymbol{x})=0$.
It should be noted that the computation of the surface integral $\eta_{ki}$ via Monte Carlo simulation is trivial, as generating samples that fall into the mutual safe domain of all other component limit-state functions poses no particular difficulty. However, if the number of the surface integrals $\eta_{ki}\;(i=1,2,\dots,n;k=1,2,\dots,m)$ is large, the computational cost for the sensitivity of system failure probability shown in Eq.(29) could still be significant. To reduce the computational effort, an importance sampling strategy that has been developed in the domain decomposition method katafygiotis2006domain for first-passage dynamic reliability analysis is introduced herein. The sum in Eq.(29) can be rewritten as
$\frac{\partial\mathbb{P}}{\partial\theta}=mn\sum_{k=1}^{m}\sum_{i=1}^{n}\eta_{ki}f(k,i)$ (35)
where $f(k,i)=1/(mn)$ defines a discrete uniform probability mass function (PMF) over the index pairs $(k,i)\;(i=1,2,\dots,n;k=1,2,\dots,m)$.
Introducing an importance sampling PMF $h(k,i)$ to Eq.(35), one yields
$\frac{\partial\mathbb{P}}{\partial\theta}=mn\sum_{k=1}^{m}\sum_{i=1}^{n}\eta_{ki}\frac{f(k,i)}{h(k,i)}h(k,i)=\sum_{k=1}^{m}\sum_{i=1}^{n}\eta_{ki}\frac{1}{h(k,i)}h(k,i)\approx\frac{1}{N}\sum_{j=1}^{N}\eta_{K_{j}I_{j}}\frac{1}{h(K_{j},I_{j})}$ (36)
where $(K_{j},I_{j})\;(j=1,2,\dots,N)$ are the index pairs sampled from $h(k,i)$, with $N$ being the number of samples; and the surface integral $\eta_{K_{j}I_{j}}$ can be estimated by Eq.(34).
There exists a theoretically optimal importance sampling PMF that results in a zero-variance estimate, i.e.
$h^{\star}(k,i)=\frac{\eta_{ki}}{\sum_{j=1}^{m}\sum_{s=1}^{n}\eta_{js}}$ (37)
However, direct use of this optimal importance sampling PMF is not possible, since the denominator $\sum_{j=1}^{m}\sum_{s=1}^{n}\eta_{js}$ is exactly the sensitivity of system failure probability $\partial\mathbb{P}/\partial\theta$ to be estimated.
A good choice of the importance sampling PMF is to construct it based on the closed-form solutions of the component failure probabilities, i.e.
$h(k,i)=\frac{\mathbb{P}_{ki}}{\sum_{j=1}^{m}\sum_{s=1}^{n}\mathbb{P}_{js}}$ (38)
The rationale of this importance sampling PMF is that a higher component failure probability $\mathbb{P}_{ki}$ typically indicates a larger measure of the constrained hyperplane $g_{ki}(\boldsymbol{x})\mathbb{I}_{ki}(\boldsymbol{x})=0$. Consequently, the corresponding surface integral $\eta_{ki}$ contributes more significantly to the sensitivity of system failure probability.
Finally, the sampling scheme described by Eqs.(34) and (36) can be combined into a single Monte Carlo estimator as follows:
$\frac{\partial\mathbb{P}}{\partial\theta}\approx\frac{1}{N}\sum_{j=1}^{N}\frac{1}{h(K_{j},I_{j})}\frac{\boldsymbol{b}_{I_{j}}^{s_{K_{j}}}\boldsymbol{x}_{j}}{\left\|\boldsymbol{a}_{I_{j}}^{s_{K_{j}}}\right\|}\mathbb{I}_{K_{j}I_{j}}(\boldsymbol{x}_{j})\varphi(\beta_{K_{j}I_{j}})$ (39)
where $N$ is the number of samples; $(K_{j},I_{j})$ is the index pair sampled from the importance sampling PMF $h(k,i)$ in Eq.(38); and $\boldsymbol{x}_{j}$ is sampled from the hyperplane $g_{K_{j}I_{j}}(\boldsymbol{x})=0$.
It should be noted that evaluating the indicator function $\mathbb{I}_{ki}(\boldsymbol{x})$ in Eq.(33) is equivalent to calling the system limit-state function $G(\boldsymbol{x})$ in Eq.(22), as both operations necessitate the computation of all component limit-state functions. Apart from evaluating the indicator function $\mathbb{I}_{ki}(\boldsymbol{x})$, the computational cost of the other operations in Eq.(39) is negligible. Therefore, the overall computational cost of the estimator in Eq.(39) amounts to $N$ evaluations of the system limit-state function $G(\boldsymbol{x})$.
Furthermore, it can be seen from Eq.(39) that the results of the indicator functions can be reused across different design parameters, meaning that the computational cost does not increase as the number of design parameters increases. This feature is particularly advantageous when sensitivities with respect to a large number of design parameters are required, as is commonly the case in reliability-based topology optimization.
For clarity, a computational procedure of the proposed surface decomposition method for the sensitivity analysis of first-passage dynamic reliability of linear systems is given as follows:
i) Perform a single impulse response time-history analysis of the linear system, and obtain the coefficients $a^{s_{k}}_{i,j}\;(1\leq j\leq i\leq n)$ for all the responses of interest $s_{k}(\boldsymbol{X},t)\;(k=1,2,\dots,m)$. Construct $\boldsymbol{a}_{i}^{s_{k}}\;(i=1,2,\dots,n;k=1,2,\dots,m)$ by Eq.(7).
ii) For each design parameter $\theta$, perform a single impulse response sensitivity time-history analysis and obtain the coefficients $b^{s_{k}}_{i,j}\;(1\leq j\leq i\leq n)$ for the response sensitivities $\partial s_{k}(\boldsymbol{X},t)/\partial\theta\;(k=1,2,\dots,m)$. Construct $\boldsymbol{b}_{i}^{s_{k}}\;(i=1,2,\dots,n;k=1,2,\dots,m)$ by Eq.(11).
iii) Compute the failure probabilities $\mathbb{P}_{ki}$, reliability indexes $\beta_{ki}$ and unit vectors $\boldsymbol{u}_{ki}$ by Eqs.(17), (18) and (20), respectively, for all component limit-state functions $g_{ki}(\boldsymbol{x})\;(i=1,2,\dots,n;k=1,2,\dots,m)$.
iv) Construct the importance sampling PMF $h(k,i)$ by Eq.(38).
v) Perform Algorithm 1 and obtain the estimate of the sensitivity of system failure probability $\partial\mathbb{P}/\partial\theta$ for each design parameter $\theta$.
Input: $\boldsymbol{a}_{i}^{s_{k}}$, $\boldsymbol{b}_{i}^{s_{k}}$ for each $\theta$, $\mathbb{P}_{ki}$, $\beta_{ki}$, $\boldsymbol{u}_{ki}$, $h(k,i)$, $tol=0.1$, $N_{\max}=10^{4}$
Output: $\partial\mathbb{P}/\partial\theta$ for each $\theta$, NN
1exfor $j\leftarrow 1$ to $N_{\max}$ do
   Generate index pair $(K_{j},I_{j})\sim h(k,i)$;
   Generate sample $\boldsymbol{x}_{j}\sim f_{\boldsymbol{X}}(\boldsymbol{x})$ and compute:
  $\boldsymbol{x}_{j}\leftarrow\beta_{K_{j}I_{j}}\boldsymbol{u}_{K_{j}I_{j}}+\big(\boldsymbol{x}_{j}-(\boldsymbol{x}_{j}^{\mathrm{T}}\boldsymbol{u}_{K_{j}I_{j}})\boldsymbol{u}_{K_{j}I_{j}}\big);$
   Compute indicator function $\mathbb{I}_{K_{j}I_{j}}(\boldsymbol{x}_{j})$;
   For each $\theta$, compute:
  $\mathrm{temp}_{j}[\theta]\leftarrow\dfrac{1}{h(K_{j},I_{j})}\dfrac{\boldsymbol{b}_{I_{j}}^{s_{K_{j}}}\boldsymbol{x}_{j}}{\left\|\boldsymbol{a}_{I_{j}}^{s_{K_{j}}}\right\|}\mathbb{I}_{K_{j}I_{j}}(\boldsymbol{x}_{j})\varphi(\beta_{K_{j}I_{j}}),$
  $\mu_{j}[\theta]\leftarrow\dfrac{1}{j}\sum_{s=1}^{j}\mathrm{temp}_{s}[\theta],$
  $\delta_{j}[\theta]\leftarrow\dfrac{\sqrt{\dfrac{1}{j-1}\sum_{s=1}^{j}(\mathrm{temp}_{s}[\theta]-\mu_{j}[\theta])^{2}}}{\left|\mu_{j}[\theta]\right|\sqrt{j}};$
  Break the loop when $\delta_{j}[\theta]<tol$ for each $\theta$;
return $\partial\mathbb{P}/\partial\theta=\mu_{j}[\theta]\;\mathrm{for\;each}\;\theta$, $N=j$
Algorithm 1 Proposed Algorithm for Surface Decomposition Method
4 Numerical examples
In this section, two numerical examples, including an oscillator and a shear-type structure, are presented to demonstrate the effectiveness of the proposed surface decomposition method (SDM) for sensitivity analysis of first-passage dynamic reliability of linear systems. The first-passage dynamic reliability analysis of the linear system is also conducted by the efficient importance sampling (ISEE) au2001first . The finite difference method in conjunction with importance sampling, which can be denoted as FDM-IS, is adopted to obtain the reference solutions to the sensitivities of failure probabilities. The basic idea of ISEE and FDM-IS are presented in A and B, respectively.
4.1 An oscillator
Consider the following equation of motion for a linear oscillator as
$\ddot{u}(\boldsymbol{X},t)+2\omega_{n}\zeta_{n}\dot{u}(\boldsymbol{X},t)+\omega^{2}_{n}u(\boldsymbol{X},t)=-\ddot{u}_{g}(\boldsymbol{X},t)$ (40)
where $\omega_{n}=4\pi\mathrm{rad/s}$ and $\zeta_{n}=0.05$ are the circular frequency and damping ratio of the linear oscillator, respectively; $\ddot{u}(\boldsymbol{X},t)$, $\dot{u}(\boldsymbol{X},t)$ and $u(\boldsymbol{X},t)$ are the acceleration, velocity and displacement of the linear oscillator, respectively; and $\ddot{u}_{g}(\boldsymbol{X},t)$ is the random seismic excitation assumed to be a stationary Gaussian white noise described by a power spectrum $S=5.5\times 10^{-4}\mathrm{m}^{2}/\mathrm{s}^{3}$. For spectral representation method, the upper and lower cutoff circular frequencies are $\omega_{\max}=25\pi$ and $\omega_{\min}=0$, respectively, while the number of circular frequency intervals is set as $q=500$. This leads to a 1000-dimensional standard Gaussian random vector $\boldsymbol{X}$.
The failure event $F$ is defined as that the peak absolute displacement of the oscillator exceeds a prescribed symmetric threshold $c$ within a given time duration $T=20\mathrm{s}$, which is expressed as
$F=\left\{c-\max_{t\in[0,T]}\left|u(\boldsymbol{X},t)\right|\leq 0\right\}$ (41)
The time step is set as $\Delta t=0.02\mathrm{s}$, resulting in a total of $n=T/\Delta t=1000$ component failure events.
The circular frequency $\omega_{n}$ and damping ratio $\zeta_{n}$ of the oscillator are selected as the design parameters. To investigate the effect of the failure probability magnitude on the efficiency of the proposed method for sensitivity analysis of dynamic reliability, four cases are considered for the symmetric threshold, i.e., $c=0.013\mathrm{m}$, $c=0.016\mathrm{m}$, $c=0.018\mathrm{m}$, and $c=0.020\mathrm{m}$. The results of failure probabilities for these cases, obtained from ISEE with a target coefficient of variation (COV) of 0.1, are presented in Table 1. The corresponding sensitivities of the failure probability with respect to $\omega_{n}$ and $\zeta_{n}$, computed by the proposed SDM with a target COV of 0.1, are shown in Table 2, together with the reference solutions obtained from FDM-IS using a 0.1% perturbation in the design parameters and a target COV of 0.02. The convergence histories of the sensitivities of failure probability, evaluated by SDM for all cases, are depicted in Figure 2.
Table 1: Results of failure probabilities for Example 1.
Case $c$ (m) $\mathbb{P}$ Function Evaluations COV
1 0.013 $3.06\times 10^{-3}$ 35 0.099
2 0.016 $1.92\times 10^{-5}$ 30 0.097
3 0.018 $3.85\times 10^{-7}$ 23 0.098
4 0.020 $4.01\times 10^{-9}$ 20 0.096
Table 2: Results of sensitivities of failure probability with respect to $\omega_{n}$ and $\zeta_{n}$ for Example 1.
Case $c$ (m) Method $\partial\mathbb{P}/\partial\omega_{n}$ COV $\partial\mathbb{P}/\partial\zeta_{n}$ COV Function Evaluations
1 0.013 SDM $7.37\times 10^{-3}$ 0.100 $5.95\times 10^{-1}$ 0.046 839
FDM-IS $7.21\times 10^{-3}$ 0.020 $6.11\times 10^{-1}$ 0.020 $6.17\times 10^{5}/1.28\times 10^{6}$
2 0.016 SDM $6.86\times 10^{-5}$ 0.100 $5.62\times 10^{-3}$ 0.054 437
FDM-IS $6.90\times 10^{-5}$ 0.020 $5.71\times 10^{-3}$ 0.020 $2.96\times 10^{5}/6.60\times 10^{5}$
3 0.018 SDM $1.73\times 10^{-6}$ 0.100 $1.44\times 10^{-4}$ 0.050 346
FDM-IS $1.60\times 10^{-6}$ 0.020 $1.42\times 10^{-4}$ 0.020 $1.98\times 10^{5}/4.65\times 10^{5}$
4 0.020 SDM $2.46\times 10^{-8}$ 0.100 $1.98\times 10^{-6}$ 0.053 267
FDM-IS $2.36\times 10^{-8}$ 0.020 $2.02\times 10^{-6}$ 0.020 $1.36\times 10^{5}/3.47\times 10^{5}$
Figure 2: Convergence histories of sensitivities of failure probability with respect to $\omega_{n}$ and $\zeta_{n}$ for Example 1. (a) Case 1: $c=0.013\mathrm{m}$; (b) Case 2: $c=0.016\mathrm{m}$; (c) Case 3: $c=0.018\mathrm{m}$; (d) Case 4: $c=0.020\mathrm{m}$. The number of function evaluations required in the proposed SDM for Case 1, Case 2, Case 3 and Case 4 are 839, 437, 346 and 267, respectively.
It can be observed from Table 1 that ISEE exhibits very high efficiency in computing the failure probabilities, with its efficiency increasing as the failure probability decreases. This phenomenon is consistent with the findings shown in au2001first , katafygiotis2006domain . Interestingly, the proposed SDM also possesses this counterintuitive characteristic in the context of sensitivity analysis, exhibiting higher efficiency as the failure probability becomes smaller, as shown in Table 2. Overall, the sensitivity results obtained from SDM are in very good agreement with the reference solutions obtained from FDM-IS, while requiring less than 1000 function evaluations, thereby demonstrating the effectiveness of the proposed method.
Case 4 is taken for further illustration. The proposed SDM requires only 267 function evaluations to compute the sensitivities of failure probability with respect to $\omega_{n}$ and $\zeta_{n}$. In fact, if the sensitivity with respect to $\zeta_{n}$ alone is of interest, fewer function evaluations are required due to its faster convergence rate, as shown in Figure 2(d), where less than 100 function evaluations are sufficient to meet the stopping criterion. For the reference method FDM-IS, $1.36\times 10^{5}$ and $3.47\times 10^{5}$ function evaluations are needed to obtain the sensitivity results with respect to $\omega_{n}$ and $\zeta_{n}$, respectively.
4.2 A shear-type structure
Consider a 20-degree-of-freedom shear-type structure, where each story is equipped with a viscoelastic damper, as illustrated in Figure 3. The mass and stiffness of each story of the structure are $m_{i}=3\times 10^{3}\mathrm{kg}$ and $k_{i}=3\times 10^{4}\mathrm{kN/m}\;(i=1,2,\dots,20)$, respectively. The Rayleigh damping model is adopted to define the damping matrix, assuming a damping ratio of 0.05 for the 1st and 20th modes of the structure. The angle between the braced arm of the viscoelastic damper and the floor is $\alpha=\arccos 0.8$. The mechanical behavior of the viscoelastic damper is characterized by the Kelvin model as follows:
$f_{ve,k}(t)=k_{ve,k}u_{ve,k}(t)+c_{ve,k}\dot{u}_{ve,k}(t)\quad(k=1,2,\dots,20)$ (42)
where $f_{ve,k}(t)$ is the damping force of the $k$th viscoelastic damper; $k_{ve,k}=3\times 10^{3}\mathrm{kN/m}$ and $c_{ve,k}=2.5\times 10^{3}\mathrm{kN\cdot s/m}\;(k=1,2,\dots,20)$ are the stiffness and damping coefficients of the $k$th viscoelastic damper, respectively; and $u_{ve,k}(t)$ and $\dot{u}_{ve,k}(t)$ are the axial nodal relative displacement and velocity between the two ends of the $k$th viscoelastic damper, respectively.
Figure 3: A 20-degree-of-freedom shear-type structure equipped with viscoelastic dampers.
The structure is subjected to a ground motion acceleration $\ddot{u}_{g}(\boldsymbol{X},t)$, which is assumed to be a uniformly modulated nonstationary Gaussian random process, i.e., $\ddot{u}_{g}(\boldsymbol{X},t)=g(t)\ddot{u}^{0}_{g}(\boldsymbol{X},t)$. $g(t)$ is the modulation function taken as
$g(t)=\begin{cases}\left(t/t_{a}\right)^{2},&0\leq t\leq t_{a},\\[6.0pt] 1,&t_{a}\leq t\leq t_{b},\\[6.0pt] e^{-\lambda(t-t_{b})},&t_{b}\leq t\leq t_{c}.\end{cases}$ (43)
where $t_{a}=8\mathrm{s}$, $t_{b}=20\mathrm{s}$, $t_{c}=30\mathrm{s}$ and $\lambda=0.1572$. $\ddot{u}^{0}_{g}(\boldsymbol{X},t)$ is the corresponding stationary Gaussian random process, with its correlation function expressed as
$R_{\ddot{u}_{g}^{0}}(\tau)=\frac{\pi S_{0}}{2}e^{-\zeta_{g}\omega_{g}|\tau|}\left(\mu_{1}\cos\omega_{d}\tau+\mu_{2}\sin\omega_{d}|\tau|\right)$ (44)
where $\tau$ denotes the time lag; $S_{0}$ is the intensity parameter directly associated with the mean peak ground acceleration; $\omega_{g}=14\mathrm{rad/s}$ and $\zeta_{g}=0.6$; and $\omega_{d}$, $\mu_{1}$ and $\mu_{2}$ are given by
$\omega_{d}=\omega_{g}\sqrt{1-\zeta_{g}^{2}},\quad\mu_{1}=\frac{\omega_{g}(1+4\zeta_{g}^{2})}{\zeta_{g}},\quad\mu_{2}=\frac{\omega_{g}(1-4\zeta_{g}^{2})}{\sqrt{1-\zeta_{g}^{2}}}.$ (45)
Consequently, the correlation function of $\ddot{u}_{g}(\boldsymbol{X},t)$ can be expressed as
$R_{\ddot{u}_{g}}(t,\tau)=g(t)g(t+\tau)R_{\ddot{u}_{g}^{0}}(\tau)$ (46)
For orthogonal decomposition method, the time step and time duration of the ground motion acceleration are taken as $\Delta t=0.02\mathrm{s}$ and $T=30\mathrm{s}$, respectively, and hence the number of time steps is $n=T/\Delta t=1500$. This results in a 1500-dimensional standard Gaussian random vector $\boldsymbol{X}$.
The failure event $F$ is defined as that the maximum absolute interstory drift among the 20 stories exceeds a given symmetric threshold $c=0.006\mathrm{m}$ within the time duration $T=30\mathrm{s}$, which can be expressed as
$F=\left\{c-\max_{i\in\left\{1,2,\dots,20\right\}}\max_{t\in[0,T]}\left|u_{i}(\boldsymbol{X},t)\right|\leq 0\right\}$ (47)
where $u_{i}(\boldsymbol{X},t)\;(i=1,2,\dots,20)$ is the interstory drift of the $i$th story. Since the number of time steps is $n=1500$ and the number of critical response components is $m=20$, a total of $mn=30000$ component failure events are considered in this example.
The stiffness and damping coefficients of the viscoelastic dampers, $k_{ve,k}$ and $c_{ve,k}\;(k=1,2,\dots,20)$, are selected as the design parameters. Three cases are considered for the intensity parameter, i.e., $S_{0}=0.010\mathrm{m^{2}/s^{3}}$, $S_{0}=0.008\mathrm{m^{2}/s^{3}}$ and $S_{0}=0.007\mathrm{m^{2}/s^{3}}$, which leads to different magnitudes of the failure probability. The results of failure probabilities are presented in Table 3, obtained from ISEE using 188, 142 and 95 function evaluations for Case 1, Case 2 and Case 3, respectively. The results of sensitivities of failure probability with respect to all the design parameters $k_{ve,k}$ and $c_{ve,k}\;(k=1,2,\dots,20)$ are depicted in Figure 4, which are obtained from the proposed SDM using 1663, 953 and 877 function evaluations for Case 1, Case 2 and Case 3, respectively. The target COV is set to be 0.1 for both ISEE and SDM. Consistent with the observation in Example 1, the ISEE and the proposed SDM exhibit higher computational efficiency as the failure probability decreases. It is seen from Figure 4 that the failure probability of the structure is most sensitive to the parameters of the first-story viscoelastic damper, i.e., $k_{ve,1}$ and $c_{ve,1}$, while the sensitivity computation with respect to the damping coefficient of the top-story viscoelastic damper, i.e., $c_{ve,20}$, dominates the convergence behavior of SDM.
Table 3: Results of failure probabilities for Example 2.
Case $S_{0}\;(\mathrm{m^{2}/s^{3}})$ $\mathbb{P}$ Function Evaluations COV
1 0.010 $3.83\times 10^{-3}$ 188 0.100
2 0.008 $3.63\times 10^{-4}$ 142 0.099
3 0.007 $7.04\times 10^{-5}$ 95 0.099
Figure 4: Results of sensitivities of failure probability with respect to $k_{ve,k}$ and $c_{ve,k}\;(k=1,2,\dots,20)$ for Example 2. (a) Case 1: $S_{0}=0.010\mathrm{m^{2}/s^{3}}$; (b) Case 2: $S_{0}=0.008\mathrm{m^{2}/s^{3}}$; (c) Case 3: $S_{0}=0.007\mathrm{m^{2}/s^{3}}$. The number of function evaluations required in the proposed SDM for Case 1, Case 2 and Case 3 are 1663, 953 and 877, respectively.
The sensitivities of the failure probability with respect to $k_{ve,1}$ and $c_{ve,1}$ are further investigated, as summarized in Table 4. The reference solutions obtained from FDM-IS, with a target COV of 0.1 and a 0.1% perturbation of design parameter, are also provided for comparison. The convergence histories of these sensitivities are illustrated in Figure 5. It is seen from Table 4 that the sensitivity results obtained from the proposed SDM align very well with the reference solutions. It is noted that the number of function evaluations for SDM reported in Table 4 corresponds to the sensitivity computation with respect to all 40 design parameters. In fact, if only the sensitivities with respect to $k_{ve,1}$ and $c_{ve,1}$ are of interest, even fewer function evaluations are required to attain the target COV of 0.1, which can be observed from Figure 5. Overall, the proposed SDM demonstrates excellent accuracy and efficiency for the sensitivity analysis of first-passage dynamic reliability of linear systems.
Table 4: Results of sensitivities of failure probability with respect to $k_{ve,1}$ and $c_{ve,1}$ for Example 2.
Case $S_{0}\;(\mathrm{m^{2}/s^{3}})$ Method $\partial\mathbb{P}/\partial k_{ve,1}$ COV $\partial\mathbb{P}/\partial c_{ve,1}$ COV Function Evaluations
1 0.010 SDM $1.06\times 10^{-9}$ 0.070 $3.93\times 10^{-9}$ 0.071 1663
FDM-IS $1.07\times 10^{-9}$ 0.100 $4.05\times 10^{-9}$ 0.100 $1.84\times 10^{6}/5.78\times 10^{5}$
2 0.008 SDM $1.56\times 10^{-10}$ 0.075 $5.75\times 10^{-10}$ 0.076 953
FDM-IS $1.54\times 10^{-10}$ 0.100 $5.56\times 10^{-10}$ 0.100 $1.05\times 10^{6}/3.52\times 10^{5}$
3 0.007 SDM $3.10\times 10^{-11}$ 0.077 $1.16\times 10^{-10}$ 0.078 877
FDM-IS $3.04\times 10^{-11}$ 0.100 $1.16\times 10^{-10}$ 0.100 $9.26\times 10^{5}/2.92\times 10^{5}$
Figure 5: Convergence histories of sensitivities of failure probability with respect to $k_{ve,1}$ and $c_{ve,1}$ for Example 2. (a) Case 1: $S_{0}=0.010\mathrm{m^{2}/s^{3}}$; (b) Case 2: $S_{0}=0.008\mathrm{m^{2}/s^{3}}$; (c) Case 3: $S_{0}=0.007\mathrm{m^{2}/s^{3}}$. The number of function evaluations required in the proposed SDM for Case 1, Case 2 and Case 3 are 1663, 953 and 877, respectively.
5 Conclusions
A novel surface decomposition method has been proposed for highly efficient and accurate sensitivity analysis of the first-passage dynamic reliability of linear systems subjected to Gaussian random excitations. To circumvent the direct computation of a complex integral over a non-smooth system limit-state hypersurface involved in the reliability sensitivity analysis, the proposed method decomposes this complex surface integral into a sum of simpler surface integrals over the constrained component hyperplanes. These component surface integrals can be evaluated at high efficiency for two main reasons: i) closed-form expressions are available for the component limit-state functions and their sensitivities for a linear system, which results in an explicit simulation of the component surface integral; ii) each constrained component hyperplane lies entirely within the mutual safe domain of all the other component limit-state functions, which leads to a straightforward generation of samples on the constrained component hyperplane. An effective importance sampling strategy is introduced to estimate the sum of all these component surface integrals, which is essential when a large number of component limit-state functions are involved.
A key advantage of the surface decomposition method lies in its ability to reuse function evaluation results for sensitivity analyses with respect to different design parameters, thereby substantially reducing the computational cost when a large number of parameters are considered. The effectiveness of the proposed method has been validated by two first-passage dynamic reliability problems of linear systems. It has been found that, for the numerical examples studied, only less than 2000 function evaluations are required to obtain the sensitivity estimates of the first-passage failure probability with respect to all design parameters with a target coefficient of variation of 0.1. Furthermore, similar to the efficient importance sampling and the domain decomposition method for first-passage dynamic reliability analysis, the surface decomposition method for reliability sensitivity analysis also demonstrates a higher efficiency as the failure probability becomes smaller. The proposed method exhibits great potential for reliability-based parametric and topology optimization.
6 Acknowledgements
This work was supported by the start-up fund and the seed fund for Basic Research for New Staff from the University of Hong Kong. The first author and the corresponding author gratefully acknowledge the support of the University of Hong Kong.
Appendix A Efficient importance sampling
Importance sampling rewrites the system failure probability integral in Eq.(21) as follows:
$\mathbb{P}=\int_{\mathbb{R}^{n}}\mathcal{H}(-G(\boldsymbol{x}))\frac{f_{\boldsymbol{X}}(\boldsymbol{x})}{h_{A}(\boldsymbol{x})}h_{A}(\boldsymbol{x})\,d\boldsymbol{x}$ (48)
where $h_{A}(\boldsymbol{x})$ denotes the importance sampling PDF, which should cover the entire system failure domain $\left\{G(\boldsymbol{X})\leq 0\right\}$.
A very efficient importance sampling PDF has been proposed in au2001first , which is given by
$h_{A}(\boldsymbol{x})=\sum_{k=1}^{m}\sum_{i=1}^{n}w_{ki}h_{ki}^{\star}(\boldsymbol{x})$ (49)
where
$h_{ki}^{\star}(\boldsymbol{x})=\frac{\mathcal{H}(-g_{ki}(\boldsymbol{x}))f_{\boldsymbol{X}}(\boldsymbol{x})}{\mathbb{P}_{ki}}\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (50)
is the theoretically optimal importance sampling PDF corresponding to the component limit-state function $g_{ki}(\boldsymbol{x})$; and
$w_{ki}=\frac{\mathbb{P}_{ki}}{\sum_{j=1}^{m}\sum_{s=1}^{n}\mathbb{P}_{js}}\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (51)
is the weight coefficient associated with the importance sampling PDF $h_{ki}^{\star}(\boldsymbol{x})$.
Substitution of Eqs.(49)-(51) into Eq.(48) yields the following importance sampling estimator:
$\mathbb{P}$ $=Z_{A}\int_{\mathbb{R}^{n}}\frac{\mathcal{H}(-G(\boldsymbol{x}))}{\sum_{k=1}^{m}\sum_{i=1}^{n}\mathcal{H}(-g_{ki}(\boldsymbol{x}))}h_{A}(\boldsymbol{x})\,d\boldsymbol{x}$ (52)
$\approx\frac{Z_{A}}{N_{A}}\sum_{j=1}^{N_{A}}\frac{1}{\sum_{k=1}^{m}\sum_{i=1}^{n}\mathcal{H}(-g_{ki}(\boldsymbol{x}_{j}))}$
where $Z_{A}=\sum_{j=1}^{m}\sum_{s=1}^{n}\mathbb{P}_{js}$ is a constant; and $\boldsymbol{x}_{j}\;(j=1,2,\dots,N_{A})$ are samples generated from $h_{A}(\boldsymbol{x})$ in Eq.(49), with $N_{A}$ being the number of samples.
Appendix B Finite difference method based on importance sampling
Finite difference method estimates the sensitivity of system failure probability as follows:
$\frac{\partial\mathbb{P}}{\partial\theta}\approx\frac{\mathbb{P}(\theta+\Delta\theta)-\mathbb{P}(\theta)}{\Delta\theta}$ (53)
where $\Delta\theta$ denotes a small perturbation in the design parameter $\theta$; and $\mathbb{P}(\theta)$ and $\mathbb{P}(\theta+\Delta\theta)$ denote the system failure probabilities corresponding to $\theta$ and its perturbed value $\theta+\Delta\theta$, respectively.
The main task of Eq.(53) is to estimate the change of probability due to the small perturbation $\Delta\theta$, which can be expressed as
$\mathbb{P}(\theta+\Delta\theta)-\mathbb{P}(\theta)=\int_{\mathbb{R}^{n}}\left[\mathcal{H}(-G(\boldsymbol{x};\theta+\Delta\theta))-\mathcal{H}(-G(\boldsymbol{x};\theta))\right]f_{\boldsymbol{X}}(\boldsymbol{x})\,d\boldsymbol{x}$ (54)
where $G(\boldsymbol{x};\theta)$ and $G(\boldsymbol{x};\theta+\Delta\theta)$ are the system limit-state functions corresponding to $\theta$ and its perturbed value $\theta+\Delta\theta$, respectively.
Importance sampling can also be used to estimate the probability integral in Eq.(54). Similar to Eq.(48), Eq.(54) can be rewritten as
$\mathbb{P}(\theta+\Delta\theta)-\mathbb{P}(\theta)=\int_{\mathbb{R}^{n}}\left[\mathcal{H}(-G(\boldsymbol{x};\theta+\Delta\theta))-\mathcal{H}(-G(\boldsymbol{x};\theta))\right]\frac{f_{\boldsymbol{X}}(\boldsymbol{x})}{h_{B}(\boldsymbol{x})}h_{B}(\boldsymbol{x})\,d\boldsymbol{x}$ (55)
A valid importance sampling PDF $h_{B}(\boldsymbol{x})$ should cover the incremental failure domain induced by the small perturbation $\Delta\theta$.
Motivated by the construction of importance sampling PDF $h_{A}(\boldsymbol{x})$ shown in Eq.(49), a valid importance sampling PDF $h_{B}(\boldsymbol{x})$ for Eq.(55) can be constructed as
$h_{B}(\boldsymbol{x})=\sum_{k=1}^{m}\sum_{i=1}^{n}w_{ki}(\theta)h_{ki}^{\star}(\boldsymbol{x};\theta)+\sum_{k=1}^{m}\sum_{i=1}^{n}w_{ki}(\theta+\Delta\theta)h_{ki}^{\star}(\boldsymbol{x};\theta+\Delta\theta)$ (56)
where
$h_{ki}^{\star}(\boldsymbol{x};\theta)=\frac{\mathcal{H}(-g_{ki}(\boldsymbol{x};\theta))f_{\boldsymbol{X}}(\boldsymbol{x})}{\mathbb{P}_{ki}(\theta)}\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (57)
and
$h_{ki}^{\star}(\boldsymbol{x};\theta+\Delta\theta)=\frac{\mathcal{H}(-g_{ki}(\boldsymbol{x};\theta+\Delta\theta))f_{\boldsymbol{X}}(\boldsymbol{x})}{\mathbb{P}_{ki}(\theta+\Delta\theta)}\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (58)
are the theoretically optimal importance sampling PDFs corresponding to the component limit-state functions $g_{ki}(\boldsymbol{x};\theta)$ and $g_{ki}(\boldsymbol{x};\theta+\Delta\theta)$, respectively; $\mathbb{P}_{ki}(\theta)$ and $\mathbb{P}_{ki}(\theta+\Delta\theta)$ are the corresponding component failure probabilities with analytical solutions available; and
$w_{ki}(\theta)=\frac{\mathbb{P}_{ki}(\theta)}{\sum_{j=1}^{m}\sum_{s=1}^{n}\left[\mathbb{P}_{js}(\theta)+\mathbb{P}_{js}(\theta+\Delta\theta)\right]}\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (59)
and
$w_{ki}(\theta+\Delta\theta)=\frac{\mathbb{P}_{ki}(\theta+\Delta\theta)}{\sum_{j=1}^{m}\sum_{s=1}^{n}\left[\mathbb{P}_{js}(\theta)+\mathbb{P}_{js}(\theta+\Delta\theta)\right]}\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (60)
are the weight coefficients associated with the importance sampling PDFs $h_{ki}^{\star}(\boldsymbol{x};\theta)$ and $h_{ki}^{\star}(\boldsymbol{x};\theta+\Delta\theta)$, respectively.
Substituting Eqs.(56)-(60) into Eq.(55), one can obtain the importance sampling estimator as follows:
$\mathbb{P}(\theta+\Delta\theta)-\mathbb{P}(\theta)$ $=Z_{B}\int_{\mathbb{R}^{n}}\frac{\mathcal{H}(-G(\boldsymbol{x};\theta+\Delta\theta))-\mathcal{H}(-G(\boldsymbol{x};\theta))}{\sum_{k=1}^{m}\sum_{i=1}^{n}\left[\mathcal{H}(-g_{ki}(\boldsymbol{x};\theta))+\mathcal{H}(-g_{ki}(\boldsymbol{x};\theta+\Delta\theta))\right]}h_{B}(\boldsymbol{x})\,d\boldsymbol{x}$ (61)
$\approx\frac{Z_{B}}{N_{B}}\sum_{j=1}^{N_{B}}\frac{\mathcal{H}(-G(\boldsymbol{x}_{j};\theta+\Delta\theta))-\mathcal{H}(-G(\boldsymbol{x}_{j};\theta))}{\sum_{k=1}^{m}\sum_{i=1}^{n}\left[\mathcal{H}(-g_{ki}(\boldsymbol{x}_{j};\theta))+\mathcal{H}(-g_{ki}(\boldsymbol{x}_{j};\theta+\Delta\theta))\right]}$
where $Z_{B}=\sum_{j=1}^{m}\sum_{s=1}^{n}\left[\mathbb{P}_{js}(\theta)+\mathbb{P}_{js}(\theta+\Delta\theta)\right]$ is a constant; and $\boldsymbol{x}_{j}\;(j=1,2,\dots,N_{B})$ are samples generated from $h_{B}(\boldsymbol{x})$ in Eq.(56), with $N_{B}$ being the number of samples.
Note that the support of the importance sampling PDF $h_{B}(\boldsymbol{x})$ is exactly the union of the two failure domains, $\left\{G(\boldsymbol{X};\theta)\leq 0\right\}$ and $\left\{G(\boldsymbol{X};\theta+\Delta\theta)\leq 0\right\}$, which sufficiently covers the incremental failure domain induced by the small perturbation $\Delta\theta$. Although the sampling region has been significantly concentrated compared with the original sampling PDF $f_{\boldsymbol{X}}(\boldsymbol{x})$, the efficiency of $h_{B}(\boldsymbol{x})$ remains limited, since the target sampling region is typically very small depending on the magnitude of the perturbation $\Delta\theta$. Nevertheless, the present finite difference method in conjunction with importance sampling provides a theoretically correct and practically feasible approach for evaluating the sensitivity of the failure probability, whereas using the original sampling PDF would be nearly impractical.